A gas sample with a mass of 12.8 g exerts a pressure of 1.2 atm at 15ºC and a volume of 9.43 L. What is the molar mass of the gas?


1
Expert's answer
2021-04-22T05:48:48-0400

PV=nRT

n=PV/RT

When the pressure is in atmospheres R=0.0821

n=(1.2x9.43)/(0.0821x288)=0.4786moles

molar mass of the gas=mass of the gas sample/number of moles

12.8/0.4786=26.74g/mol